(x – 2) + (x – 3) + (x – 9) = 0
(x – 2) + (x – 3) + (x – 9) = 0
(x – 2) + (x – 3) + (x – 9) = 0
(x – 2) + (x – 3) + (x – 9) = 0
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(x – 2) + (x – 3) + (x – 9) = 0

x – 2 + x – 3 + x – 9 = 0

3x – 2 – 3 – 9 = 0

3x – 14 = 0

x = 14/3

Thanks